Hi..I try to solve this..but I got lost
At Fictional University of Technology, 65% of incoming freshmen students are male and 35% are female. Recent records indicate that 80% of entering male students will graduate with a B.Sc. degree, while 85% the female students will obtain a B.Sc. degree. If an incoming freshman student is selected at random, find
a. P(student will graduate | students is female)
b. P(student will graduate and student is female)
c. P(student will graduate)
d. If you answers in part “a” and “b” are the same, explain why. If they are different, explain why.
thx

Answer by stanbon(75887)

At Fictional University of Technology, 65% of incoming freshmen students are male and 35% are female. Recent records indicate that 80% of entering male students will graduate with a B.Sc. degree, while 85% the female students will obtain a B.Sc. degree. If an incoming freshman student is selected at random, find
a. P(student will graduate | students is female) = 0.85
-----------------------------
b. P(student will graduate AND student is female)
= P(female)*P(grad|female) = 0.35*0.85
--------------------------------------------
c. P(student will graduate) = 0.8*0.65+0.85*0.35
-------------------------------------------------------
d. If you answers in part “a” and “b” are the same, explain why. If they are different, explain why.
I'll leave that to you.
==========================
Cheers,
Stan H.
==========================